Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  RH-ISAC, Page 105 of 140  Not logged in ELOG logo
ID Date Author Category Type Module Target/Number Subject Text Attachments
  722   Tuesday, August 05, 2014, 11:53 David WangITERepair  Reconnected ITE entrance moduleITE entrance module is reconnected in ITE.   
  721   Tuesday, August 05, 2014, 10:40 Travis CaveITEStandard Operation  module moveITE Entrance module has been moved back
to the east station.
  
  720   Thursday, July 31, 2014, 08:08 chad fisherSouth Hot-CellStandard OperationTM3SiC#30SiC#30 Removed and inpectSiC#30 has been removed from TM3. A quick
exam of TM3 source tray did not  yield
any obvious signs of damage.
 TM320140731_01.jpgTM320140731_02.jpgTM320140731_03.jpgTM320140731_04.jpg 
  719   Wednesday, July 30, 2014, 11:51 Travis CaveConditioning StationMaintenance  entrance module moveITE Entrance module has been moved from
the east station to the conditioning station.
  
  718   Monday, July 28, 2014, 17:26 Ron KuramotoITWRepair  ITW:IV5 Air leak repairAir leak
was found at the Solenoid control valve located
in the Pre-Separator & DB0 room.  The
 DSCN0087.JPG20140728_142855_-_small.jpg20140728_143202_-_small.jpg 
  717   Monday, July 28, 2014, 14:45 David WangConditioning StationDevelopmentTM2 turned off TM2 and CS turbo pumps.I turned off TM2 and CS turbo-pumps. TM2
will be moved from CS to silo tomorrow morning.
  
  716   Monday, July 28, 2014, 14:37 David WangConditioning StationDevelopmentTM2 helium pressure test on TM2 tube heater +(D)and -(C)water lines.
    TM2 installed in Conditioning
Station
    Varian Leak detector
connected directly to Target Module 2   
 7x 
  715   Monday, July 28, 2014, 09:29 Travis CaveConditioning StationStandard OperationTM2no targetmodule moveTM#2 has been moved from the south hot
cell to the conditioning station.
  
  714   Thursday, July 24, 2014, 18:21 Grant MinorSouth Hot-CellRepairTM2No target / no source trayLeak check on TM2 ionizer tube cooling lines C and D - module sideTarget Ionizer tube cooling lines "C"
and "D" in Target Module 2 were
both leak checked at the Hot Cell. 
 ITA6021_Rev_14_screencapture_cooling_lines_24July2014.jpgTM2_hot_cell_leak_check_tube_heater_C_D_24July2014_small2.jpgTM2_hot_cell_leak_check_no_source_tray_24july2014.JPG 
  713   Thursday, July 24, 2014, 09:10 Travis CaveSouth Hot-CellStandard OperationTM2no targetmodule moveTM#2 has been moved to the hot cell from
the north east silo.
  
  712   Wednesday, July 23, 2014, 13:36 David WangITWStandard OperationTM1 TM1 in ITW.  TM1 has been connected in ITW .
The new roughing pump replaced on last  Friday
works good. I started all turbo pumps in
  
  711   Tuesday, July 22, 2014, 07:32 David WangConditioning StationStandard OperationTM1 Stop TM1 and CS turbo pumps and ruff pump.Stop TM1 and CS turbo pumps and ruff pump.
TM1 is disconnect at CS. It will be vented
tomorrow morning and moved to ITW.
  
  710   Monday, July 21, 2014, 13:11 David WangConditioning StationStandard OperationTM1 helium pressure test on TM1 HS water line at CS.The TM1 HS line is  pressure tested
this morning with 70 psi helium  at
CS. No response has been found on leak detector.
  
  709   Thursday, July 17, 2014, 11:49 David WangConditioning StationStandard OperationTM1 TM1 and CS turbo pumps have been startedI connected TM1 at CS. All turbo pumps
have been started. The pumping down is 
good so far.
  
  708   Thursday, July 17, 2014, 10:13 David WangConditioning StationStandard OperationTM1 Leak check on TM1 HS water line with new target installedChad and I leak checked TM1 HS waterline
with new target installed at SHC. The LD
base LR:0.0xE-9atm.cc/sec, base pressure:
  
  707   Thursday, July 17, 2014, 09:29 Travis CaveSouth Hot-CellStandard OperationTM1Ta#43module moveTM#1 with Ta#43 has been moved from the
south hot cell to conditioning station.
  
  706   Wednesday, July 16, 2014, 14:28 Travis CaveSouth Hot-CellStandard OperationTM1Ta#43Electrical checkTM#1 with Ta#43 electrical check see attached
PDF.
 TM1_Ta43_elec_chk_shc.PDF 
  705   Wednesday, July 16, 2014, 13:59 chad fisherSouth Hot-CellStandard OperationTM3SiC#30Sic#30 installJust some notes on the install of SiC#30
on TM3 last month...It was installed with
all new fasteners and torque values provided
  
  704   Wednesday, July 16, 2014, 13:56 chad fisherSouth Hot-CellStandard OperationTM1UCx#9UCx#9 off Ta onThe UCx#9 target has been revomed from
the target module...leg A was broken as suspected.
Ta#42 has been installed. High current
  
  703   Wednesday, July 16, 2014, 13:02 Travis CaveSouth Hot-CellStandard OperationTM1UC#9Electrical checkElectrical check of the TM#1 with UC#9
post beam at the south hot cell.
 TM1_UC9_Elecchk_postbeam.PDF 
ELOG V2.9.2-2455